Probability that a plant will live is 3/4 and the probability that another plant lives is 1/3. The probability that only one of them lives is?

OPTIONS
  • 7/12
  • 1/4
  • 1/6
  • 1/5
  • NONE OF THE ABOVE
​PLEASE EXPLAIN WITH FULL STEPS

Grade:12th Pass

1 Answers

Sandeep Bhamoo
4 years ago
Answer is first option 7/12. Since only one should live then there are 2 cases -first dies,second lives P=(1/4)(1/3) = 1/12. Second dies and first lives P= (3/4)(2/3) = 6/12. Hebce total probability =6/12 +1/12. =7/12. Please do approve my answer if correct